# Federated Prototype-based Contrastive Learning for Privacy-Preserving Cross-domain Recommendation

### v1

@@ IV-B Experimental Results and Analysis
We evaluate the performance of FedPCL-CDR and the baselines using commonly used evaluation metrics w.r.t. HR@10 and NDCG@10. From the experimental results in Table I, We can observe that:
@@
  • Our model, FedPCL-CDR, surpasses other baselines, showcasing its capability to achieve satisfactory performance while also safeguarding user privacy. Specifically, FedPCL-CDR outperforms the top-performing CDR baseline by an average of 7.38% in HR@10 and 4.86% in NDCG@10 across all tasks. This improvement can be attributed to the following reasons: (1) FedPCL-CDR efficiently utilizes non-overlapping user data to transfer cross-domain knowledge, which is particularly beneficial in scenarios with sparse overlapping users, such as Phone&Sport. (2) By constructing prototype-based CL tasks, FedPCL-CDR achieves more effective knowledge transfer.
@@
  • FedPCL-CDR outperforms single-domain federated methods such as FedNCF. This demonstrates the significant role of cross-domain knowledge in enhancing recommendation performance within the FL framework. In addition, FedPCL-CDR performs better than PriCDR and P2FCDR, indicating that methods leveraging prototypes and FL for user privacy protection surpass those utilizing differential privacy technology. Moreover, FedPCL-CDR exceeds the performance of GA-DTCDR, which depends on fully overlapping users for knowledge transfer. This indicates that effectively utilizing non-overlapping user information can improve model performance. Finally, although our method and CL-DTCDR both use CL to transfer knowledge, FedPCL-CDR still performs better than CL-DTCDR, showing that our method not only protects user privacy but also improves model performance.
@@
  • GNN-based methods outperform non-graph methods, such as LightGCN vs NeuMF. This demonstrates that incorporating high-order neighbor information can improve model accuracy.
@@
To validate the effectiveness of each component in FedPCL-CDR, we conducted ablation experiments. We created two variants: (1) w/o loc-proto: we eliminate the local prototype-based CL loss. (2) w/o glob-proto: we remove the global prototype-based CL loss. The results of the ablation studies are shown in Table II. We can observe that: (1) The inferior performance of models w/o loc-proto and w/o glob-proto demonstrates the significant contributions of local and global prototype-based CL to the outcome. (2) In general, w/o loc-proto contributes more, which shows that local prototype-based CL plays an important role in improving model performance. In conclusion, each component in FedPCL-CDR plays a crucial role, demonstrating the rationality and effectiveness of our design.
TABLE II: Ablation studies on FedPCL-CDR.
Task Domain Metric w/o loc-proto w/o glob-proto FedPCL-CDR
Task 1 Phone HR@10 0.5796 0.5870 0.5933
NDCG@10 0.3853 0.3862 0.3949
Sport HR@10 0.6610 0.6721 0.6766
NDCG@10 0.4468 0.4495 0.4534
Task 2 Elec HR@10 0.5980 0.6011 0.6087
NDCG@10 0.3908 0.3931 0.4018
Cloth HR@10 0.4954 0.5006 0.5018
NDCG@10 0.3123 0.3137 0.3166
Phone HR@10 0.5863 0.5811 0.5970
NDCG@10 0.3733 0.3743 0.3844
Task 3 Book HR@10 0.4915 0.4967 0.5009
NDCG@10 0.2896 0.2876 0.2980
Movie HR@10 0.4608 0.4595 0.4621
NDCG@10 0.2352 0.2444 0.2450
Task 4 Book HR@10 0.4688 0.4677 0.4724
NDCG@10 0.2727 0.2769 0.2865
Movie HR@10 0.4506 0.4514 0.4601
NDCG@10 0.2291 0.2285 0.2444
Music HR@10 0.4532 0.4388 0.4745
NDCG@10 0.2619 0.2514 0.2691
IV-B3 Performance for different proportions of overlapping users.
@@
To assess FedPCL-CDR’s capability in addressing sparse overlapping users within CDR, we manipulate the overlapping ratio specifically for Task 1 across different settings. These varying ratios signify different levels of commonality, where a higher ratio indicates a greater number of overlapping users across domains. For instance, in Task 1 with the “Phone-Sport” dataset and an overlapping ratio of 30%, the number of overlapping users is computed as 655 * 30% = 196. Due to space limitations, we report results about several representative CDR methods. The corresponding results with different overlapping ratios are shown in Table III. As the overlapping ratio increases, the performance of all models demonstrates improvement. This is intuitively sensible, as a higher overlapping ratio implies a greater number of shared users, facilitating a more straightforward transfer of cross-domain knowledge. The performance of GA-DTCDR and PriCDR shows significant fluctuations, primarily because they heavily rely on fully overlapping users. In contrast, both NMCDR, CL-DTCDR and our method, FedPCL-CDR, demonstrate relatively minor changes, indicating that effectively transferring knowledge across non-overlapping users can enhance performance and ensure model stability.
@@
TABLE III: Experimental results on Task 1 with different overlapping user ratios.
@@
Domain overlapping user ratio Metric GA-DTCDR NMCDR PriCDR CL -DTCDR PedPCL -CDR
Phone 30% HR 0.2042 0.3649 0.2041 0.4954 0.5831
NDCG 0.1022 0.1994 0.1053 0.2685 0.3852
50% HR 0.2691 0.3762 0.2110 0.5047 0.5861
NDCG 0.1426 0.2078 0.1185 0.2736 0.3886
70% HR 0.2838 0.3733 0.2449 0.5124 0.5909
NDCG 0.1596 0.2037 0.1464 0.2731 0.3825
100% HR 0.3809 0.3841 0.3753 0.5277 0.5933
NDCG 0.2285 0.2349 0.2259 0.3193 0.3949
Sport 30% HR 0.2520 0.3746 0.1429 0.5292 0.6664
NDCG 0.1413 0.1951 0.0743 0.3173 0.4453
50% HR 0.2910 0.3863 0.1865 0.5381 0.6696
NDCG 0.1713 0.2026 0.1002 0.3203 0.4475
70% HR 0.3144 0.3908 0.2271 0.5493 0.6706
NDCG 0.1983 0.2128 0.1286 0.3302 0.4508
100% HR 0.3749 0.3914 0.4087 0.5660 0.6766
NDCG 0.2056 0.2166 0.2389 0.3344 0.4534
IV-B4 Empirical Study of Privacy.
@@
In this subsection, we demonstrate the privacy-preserving capabilities of our FedPCL-CDR model by simulating an attack in which an attacker attempts to reconstruct the original user embeddings from intercepted prototypes. We analyze the difficulty of this reconstruction and evaluate the effectiveness of our privacy-preserving mechanisms. We assume that an attacker intercepts the local prototypes during the client-server communication. The attacker employs a deep neural network model to infer the original user embeddings from these intercepted prototypes, training the model to minimize the difference between the reconstructed and actual embeddings. We use Mean Squared Error (MSE) as an evaluation metric to measure the accuracy of the reconstructed embeddings compared to the original ones. We conducted experiments on Tasks 1 and 2 and reported the results in Table IV. These high MSE values indicate that the attacker faces significant difficulty in accurately inferring the original user embeddings from the intercepted prototypes, demonstrating the strong privacy-preserving capabilities of our model.
TABLE IV: MSE on Tasks 1 and 2.
Task Task 1 Task 2
Domain Phone Sport Elec Cloth Phone
MSE 6.79 10.22 6.80 7.74 5.20
IV-B5 Impact of Hyper-parameters.
@@
In this section, we assess the model’s performance across different configurations of two pivotal parameters: the weight parameter $\alpha$ associated with prototype-based CL losses, and the cluster number $c$. Due to space limitations, we only report HR@10 and NDCG@10 on Tasks 1 (Phone&Sport) and 3 (Movie&Music).
@@
  • Impact of $\alpha$. We employ the parameter $\alpha$ to control the degree of knowledge transfer across domains. To evaluate its impact, we conduct experiments with different $\alpha$ values, namely $[0.001,0.01,0.1,0.2]$. Figure 3 illustrates the outcomes for HR@10 and NDCG@10. We find that as the weight parameter $\alpha$ increases, the performance first rises and then decreases. FedPCL-CDR achieves optimal results when $\alpha$ is set to 0.01.
    (a) Phone&Sport
    (b) Movie&Music
@@
  • Impact of cluster number $c$. The number of clusters significantly influences the generalization of prototypes, thereby affecting the learning of user preferences. As shown in Figure 4, our model reaches its peak performance when the number of clusters is set to 10. With the increase in the number of clusters, the HR@10 and NDCG@10 metrics initially rise, reaching a maximum at $c=10$, and subsequently decline. This trend can be attributed to the fact that an excessive number of clusters results in overly specific prototypes, which lack generalization capabilities and lead to suboptimal knowledge transfer.
    (a) Phone&Sport
    (b) Movie&Music
    Figure 4: Performance of different cluster number c.
V Conclusion and Future work
In this paper, we propose a Federated User Preference Modeling framework (FUPM) for PPCDR to address the data sparsity issue while protecting user privacy. Within FUPM, we first design a comprehensive preference exploration module to learn comprehensive user preferences from review texts and potentially positive items. We then devise a private preference transfer module to privately transfer user preferences within the FL framework. Importantly, in order to protect user privacy during cross-domain knowledge transfer, we learn local prototypes and apply LDP techniques to them before transfer. Extensive experimental results on four CDR tasks based on real-world Amazon and Douban datasets demonstrate the effectiveness of our proposed FUPM.
Our study assumes overlapping users and non-overlapping items across domains. While our method can be extended to scenarios with partially overlapping users, it is not applicable to scenarios with no user overlap or only partial item overlap. Future work includes exploring effective methods to address these challenges.

### v2

@@ 4.2.1 Performance Evaluation (RQ1)
FedPCL-CDR and baselines are evaluated using standard metrics including HR@10 and NDCG@10. From the experimental results in Table 2, We can observe that:
  • 1.
@@
    Our model, FedPCL-CDR, surpasses other baselines, showcasing its capability to achieve satisfactory performance while also safeguarding user privacy. Specifically, FedPCL-CDR outperforms the top-performing CDR baseline by an average of 2.45% in HR@10 and 1.93% in NDCG@10 across all tasks. This improvement can be attributed to the following reasons: (1) FedPCL-CDR efficiently utilizes non-overlapping user data to transfer cross-domain knowledge, which is particularly beneficial in scenarios with sparse overlapping users, such as Phone&Sport. (2) By constructing dual prototype-based CL tasks, FedPCL-CDR achieves more effective knowledge transfer.
  • 2.
@@
    FedPCL-CDR outperforms single-domain federated methods such as FedNCF. This demonstrates the significant role of cross-domain knowledge in enhancing recommendation performance within the FL framework. Furthermore, FedPCL-CDR outperforms the PPCDR methods PriCDR and P2FCDR, demonstrating that leveraging prototypes for knowledge transfer not only protects user privacy but also improves model performance. Moreover, FedPCL-CDR performs better than $P^{2}$, demonstrating that utilizing non-overlapping user information and dual-prototype CL learning paradigm can improve the recommendation performance. Finally, although our method and CL-DTCDR both use CL to transfer knowledge, FedPCL-CDR still performs better than CL-DTCDR, showing that our method not only protects user privacy but also improves model performance.
  • 3.
@@
    GNN-based methods (e.g., LightGCN) surpass non-graph approaches (e.g., NeuMF), which shows that leveraging high-order neighborhood information improves accuracy.
  • 4.
@@ 4.2.2 Ablation Studies (RQ2)
To evaluate FedPCL-CDR’s components, we created three variants: (1) w/o loc-proto: we eliminate the local prototype-based CL loss. (2) w/o glob-proto: we remove the global prototype-based CL loss. (3) w/o LDP: We remove the LDP mechanism during FedPCL-CDR training, enabling direct prototype transfer across domains without noise injection. Table 3 presents the results. We can observe that: (1) The performance drop in models w/o loc-proto and w/o glob-proto highlights the crucial role of both local and global prototype-based CL in achieving superior performance. (2) In general, w/o loc-proto contributes more, which shows that local prototype-based CL plays an important role in improving model performance. (3) While w/o LDP yields higher accuracy by omitting noise injection, this variant increases the risk of privacy leakage. In conclusion, each component in FedPCL-CDR plays a crucial role, demonstrating the rationality and effectiveness of our design.
Table 3: Ablation studies on FedPCL-CDR.
Task Domain Metric w/o loc-proto w/o glob-proto w/o LDP FedPCL-CDR
Task 1 Phone HR 0.5478 0.5364 0.5933 0.5501
NDCG 0.3532 0.3487 0.3949 0.3541
Sport HR 0.5868 0.5938 0.6766 0.5991
NDCG 0.3760 0.3788 0.4534 0.3812
Task 2 Elec HR 0.5371 0.5248 0.6078 0.5408
NDCG 0.3413 0.3356 0.4018 0.3445
Cloth HR 0.4324 0.4455 0.5018 0.4489
NDCG 0.2641 0.2590 0.3166 0.2678
Phone HR 0.5222 0.5240 0.5970 0.5341
NDCG 0.3304 0.3312 0.3844 0.3417
Task 3 Movie HR 0.4232 0.4241 0.4379 0.4336
NDCG 0.2291 0.2293 0.2356 0.2301
Music HR 0.3203 0.3194 0.4518 0.3294
NDCG 0.1801 0.1736 0.2487 0.1842
Task 4 Book HR 0.3656 0.3662 0.4724 0.3727
NDCG 0.2171 0.2058 0.2865 0.2265
Movie HR 0.4414 0.4444 0.4610 0.4470
NDCG 0.2317 0.2283 0.2444 0.2334
Music HR 0.3401 0.3385 0.4745 0.3418
NDCG 0.1907 0.1860 0.2059 0.1935
4.2.3 Performance for different proportions of overlapping users (RQ3)
@@
To assess FedPCL-CDR’s capability in addressing sparse overlapping users within CDR, we manipulate the overlapping ratio specifically for Task 1 across different settings. These varying ratios signify different levels of commonality, where a higher ratio indicates a greater number of overlapping users across domains. For instance, in Task 1 with the “Phone-Sport” dataset and an overlapping ratio of 30%, the number of overlapping users is computed as 655 * 30% = 196. We report results about several representative CDR methods. The corresponding results with different overlapping ratios are shown in Table 4. As the overlapping ratio increases, the performance of all models demonstrates improvement. This is intuitively sensible, as a higher overlapping ratio implies a greater number of shared users, facilitating a more straightforward transfer of cross-domain knowledge. The performance of GA-DTCDR, PriCDR, and $P^{2}$ shows significant fluctuations, primarily because they rely on overlapping users to transfer knowledge. In contrast, both NMCDR, CL-DTCDR, and our method, FedPCL-CDR, demonstrate relatively minor changes, indicating that effectively transferring knowledge across non-overlapping users can enhance performance and ensure model stability.
@@
Table 4: Experimental results on Task 1 with different overlapping user ratios.
@@
Domain overlapping user ratio Metric GA-DTCDR NMCDR PriCDR CL -DTCDR $P^{2}$ PedPCL -CDR
Phone 30% HR 0.2042 0.3649 0.2041 0.4954 0.3014 0.5223
NDCG 0.1022 0.1994 0.1053 0.2685 0.1874 0.3368
50% HR 0.2691 0.3762 0.2110 0.5047 0.3204 0.5353
NDCG 0.1426 0.2078 0.1185 0.2736 0.2018 0.3402
70% HR 0.2838 0.3733 0.2449 0.5124 0.3578 0.5412
NDCG 0.1596 0.2037 0.1464 0.2731 0.2201 0.3421
Sport 30% HR 0.2520 0.3746 0.1429 0.5292 0.3604 0.5649
NDCG 0.1413 0.1951 0.0743 0.3173 0.1903 0.3557
50% HR 0.2910 0.3863 0.1865 0.5381 0.3892 0.5781
NDCG 0.1713 0.2026 0.1002 0.3203 0.2213 0.3679
70% HR 0.3144 0.3908 0.2271 0.5493 0.4110 0.5845
NDCG 0.1983 0.2128 0.1286 0.3302 0.2403 0.3746
4.2.4 Empirical Study of Privacy (RQ3)
@@
We compare the privacy-preserving capabilities of FedPCL-CDR and PPCDR baselines by simulating an attack where an attacker attempts to reconstruct original user embeddings. Assuming the attacker intercepts user features during client-server communication, they use a deep neural network to infer the original embeddings, minimizing the reconstruction error. We use Mean Squared Error (MSE) as an evaluation metric to measure the accuracy of the reconstructed embeddings compared to the original ones. A higher MSE value indicates greater resistance to original embedding reconstruction, demonstrating stronger privacy preservation. We conducted experiments on Tasks 1 and 3 and reported the results in Table 5. We can find that FedPCL-CDR obtains the optimal privacy-preserving capabilities.
Additionally, we analyze the trade-off between recommendation performance and privacy, as shown in Figure 3. A larger privacy budget $\epsilon$ results in improved performance but leads to greater privacy leakage. Notably, Notably, performance variation becomes relatively stable when $\epsilon>=4$, indicating this threshold represents an optimal privacy-accuracy trade-off. Therefore, we set $\epsilon=4$ to achieve a balance between performance and privacy.
Table 5: MSE on Tasks 1 and 3.
Task Domain Metric P2FCDR $P^{2}DTR$ FedPCL-CDR
Task 1 Phone MSE 5.12 6.74 8.72
Sport MSE 9.25 10.25 12.67
Task 3 Movie MSE 3.21 3.78 4.48
Music MSE 3.46 4.32 5.67
((a))
((b))
Figure 3: Performance of different privacy budgets on Tasks 1 and 3.
4.2.5 Impact of Hyper-parameters (RQ4)
@@
We assess FedPCL-CDR’s performance across different configurations of three pivotal parameters: the weight parameter $\alpha$ associated with prototype-based CL losses, the cluster number $K$, and the recommendation list length $N$.
  • 1.
@@
    Impact of $\alpha$. We employ the parameter $\alpha$ to control the degree of knowledge transfer across domains. To evaluate its impact, we conduct experiments with different $\alpha$ values, namely $[0.001,0.01,0.1,0.2]$. Figure 4 illustrates the outcomes for HR@10 and NDCG@10. We find that as the weight parameter $\alpha$ increases, the performance first rises and then decreases. The best performance of FedPCL-CDR is observed at $\alpha=0.01$.
    ((a))
    ((b))
@@
  • 2.
    Impact of cluster number $K$. The number of clusters significantly influences the generalization of prototypes, thereby affecting the learning of user preferences. As shown in Figure 5, FedPCL-CDR reaches its peak performance when the number of clusters is set to 10. With the increase in the number of clusters, the HR@10 and NDCG@10 metrics initially rise, reaching a maximum at $K=10$, and subsequently decline. This trend can be attributed to the fact that an excessive number of clusters results in overly specific prototypes, which lack generalization capabilities and lead to suboptimal knowledge transfer.
    ((a))
    ((b))
    Figure 5: Performance of different cluster number K.
  • 3.
    Impact of recommendation list length $N$. To assess how the length of the recommendation list influences performance, we evaluate FedPCL-CDR under different values of $N\in\{2,4,6,8,10\}$, as shown in Figure 6. The results reveal a consistent performance improvement as $N$ increases. This trend is expected, since recommending a greater number of items increases the likelihood of including relevant ones, thereby simplifying the task and boosting model effectiveness.
    ((a))
    ((b))
    Figure 6: Performance of different recommendation list length N.
5 Conclusion and Future work
In this paper, we introduce FedPCL-CDR, a federated prototype-based CL framework for PPCDR. It aims to address the sparse overlapping problem and enhance user privacy preservation. FedPCL-CDR comprises two modules: (1) a local domain learning module that privately transfers knowledge of both overlapping and non-overlapping users across domains using differential prototypes in a CL manner, and (2) a global server aggregation module that aggregates user interests from multiple domains by modeling local and global prototypes. We have conducted extensive experiments on four CDR tasks using two real-world datasets to demonstrate the effectiveness of our proposed model.
While our proposed framework, FedPCL-CDR, effectively addresses the sparse overlapping user problem while enhancing privacy protection, it has several limitations. First, although prototype transmission reduces communication costs, frequent server-client synchronization in federated learning may still introduce latency in cross-device deployments. Second, while our method can operate with very sparse overlaps, it fundamentally requires some overlapping users to select representative prototypes, making fully non-overlapping scenarios unsolved. Finally, the fixed cluster number may not adapt well to domains with highly dynamic user interests. We plan to address these limitations in future work.
Acknowledgements
This work is supported by the Australian Research Council (LP210100129).
CRediT authorship contribution statement
Li Wang: Conceptualization, Investigation, Methodology, Software, Writing - original draft, Writing - review & editing. Qiang Wu: Supervision, Writing - review & editing. Min Xu: Supervision, Writing - review & editing.